OpenStudy (anonymous):

If $535 is invested at an interest rate of 6% per year and is compounded continuously, how much will the investment be worth in 10 years? So I plugged that in just like someone showed me and I got zero??

OpenStudy (anonymous):

remember Pe^rt

OpenStudy (anonymous):

whenever they say continous compounding

OpenStudy (anonymous):

\[A=Pe^{rt}\]

OpenStudy (anonymous):

A= 535 e^(.06*10) now just evaluate this

OpenStudy (anonymous):

Okay so wait what about e?

OpenStudy (anonymous):

most calculator has 'e' button

OpenStudy (anonymous):

Okay, so once I multiply .06 *10 what do I do next? divide or multiply? Oh okay, so I look for e. alright that makes sense.

OpenStudy (anonymous):

I thought I had tos olve for e or something.

OpenStudy (anonymous):

A= 535 e^(.6)

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=535+e^%28.6%29

OpenStudy (anonymous):

Wow, thanks :). So for: A bacteria culture begins with 4 bacteria which double in size every hour. How many bacteria exist in the culture after 8 hours? I'd do A=4 e ^(.8)

OpenStudy (anonymous):

\[A=A_0 e^{rt}\] This is the formula \[A_0\] is initial state ----------------------------------------------- \[r\] is rate --------------------------------------------- \[t\] is time

OpenStudy (anonymous):

we have to find r first

OpenStudy (anonymous):

2? ^2?

OpenStudy (anonymous):

every 60 minutes?

OpenStudy (anonymous):

A bacteria culture begins with 4 bacteria which double in size every hour \[A=A_0e^{rt}\] \[8=4e^{r1}\] \[2=e^{r1}\] \[ln(2)=ln(e^r)\] \[ln(2)=r * ln (e)\] \[ln(2)=r \]

OpenStudy (anonymous):

So are they two different equations/

OpenStudy (anonymous):

actually , they are same we just use different letters

OpenStudy (anonymous):

So so i plug anything into ln(2)=r

OpenStudy (anonymous):

yes,

OpenStudy (anonymous):

Or I can just plug it into a=ae^rt

OpenStudy (anonymous):

\[A=A_0 e^{ln(2)t}\]

OpenStudy (anonymous):

Is the answer 1024?

OpenStudy (anonymous):

That's what I got.

OpenStudy (anonymous):

\[A=4e^{ln(2)8}\] Yes 1024 GJ

OpenStudy (anonymous):

Thank you soo much :)